# algorithmic

### v1

@@ Abstract
In this paper, we study a best arm identification problem with dual objects. In addition to the classic reward, each arm is associated with a cost distribution and the goal is to identify the largest reward arm using the minimum expected cost. We call it Cost Aware Best Arm Identification (CABAI), which captures the separation of testing and implementation phases in product development pipelines and models the objective shift between phases, i.e., cost for testing and reward for implementation. We first derive an theoretic lower bound for CABAI and propose an algorithm called $\mathsf{CTAS}$ to match it asymptotically. To reduce the computation of $\mathsf{CTAS}$, we further propose a low-complexity algorithm called CO, based on a square-root rule, which proves optimal in simplified two-armed models and generalizes surprisingly well in numerical experiments. Our results show (i) ignoring the heterogeneous action cost results in sub-optimality in practice, and (ii) low-complexity algorithms deliver near-optimal performance over a wide range of problems.
@@ 1 Introduction
The stochastic multi-armed bandit (MAB) (Thompson, 1933; Robbins, 1952) is a classic model which has widespread applications, from content recommendation (Kohli et al., 2013), resource allocation (Liu et al., 2020), clinical trials (Villar et al., 2015), to efficient ad placement. A multi-armed bandit problem involves an agent and an environment, which is represented by a set of actions (arms) with distinct underlying reward distributions. At each round, the agent will choose one of the arms and then obtain a random reward generated from the associated distribution. Most existing studies formulate MAB either as the best arm identification (BAI) problem (Kaufmann et al., 2016a; Garivier & Kaufmann, 2016) where the agent intends to identify the highest reward arm as quickly as possible or as a regret minimization problem (Auer, 2002; Garivier & Cappé, 2011) where the goal is to maximize the cumulative reward over a time-horizon. Both formulations have been well-developed and successful in balancing the trade-off between exploration and exploitation.
@@
Problem Algorithm Optimal? $w_{1}(t)$ $(1.5,1)$ $w_{2}(t)$ $(1,0.1)$ $w_{3}(t)$ $(0.5,0.01)$
BAI $\mathsf{TAS}$ $0.46$ $0.46$ $0.08$
CABAI $\mathsf{CTAS}$ $0.23$ $0.72$ $0.05$
@@ 1.1 Related Work
BAI with Fixed Confidence: BAI has been studied for many years and was originally proposed in (Bechhofer, 1958). In this paper, we consider a subset known as the fixed confidence setting, where the agent aims to minimize the sample complexity while ensuring the best arm is identified with probability at least $1-\delta$. Here, $\delta$ is a pre-specified confidence level, and such algorithms are called $\delta$-PAC. In (Kaufmann et al., 2016a), the authors introduce a non-asymptotic lower bound for this setting. Subsequently, they propose the Track and Stop algorithm ($\mathsf{TAS}$) that matches this lower bound asymptotically. The $\mathsf{TAS}$ algorithm has since been extended to various other settings (Jourdan et al., 2023; Garivier & Kaufmann, 2021; Kato & Ariu, 2021). Before it, researchers proposed “confidence-based” algorithms, e.g., $\mathsf{KL}$-$\mathsf{LUCB}$ (Kaufmann & Kalyanakrishnan, 2013), $\mathsf{UGapE}$ (Gabillon et al., 2012), which rely on constructing high-probability confidence intervals. They are more computationally feasible than $\mathsf{TAS}$ inspired algorithms, but with few theoretical guarantees.
BAI with Safety Constraints: A formulation similar to our paper is BAI with safety constraints (Wang et al., 2022). As a motivating example, they consider the clinical trial setting, where each drug is associated with a dosage and the dosage has an associated safety level. They attempt to identify the best drug and dosage for fixed confidence without violating the safety level. Similarly, Hou et al. (2022) attempts to identify the best arm subject to a constraint on the variance. Our formulation is distinct from them because the agent is free to perform any action.
@@ 2 Preliminaries
We study a model similar to the fixed-confidence BAI in stochastic $K$-armed bandits. We denote the set of arms as $\mathcal{A}\coloneqq\{1,2,\cdots,K\}$. Each arm $a$ is associated with a reward distribution $\boldsymbol{\nu}_{\mu}=\{\nu_{\mu_{1}},\ldots,\nu_{\mu_{K}}\}$ with expectations $\boldsymbol{\mu}\coloneqq\{\mu_{1},\mu_{2},\cdots,\mu_{K}\}$. We assume $\boldsymbol{\nu}_{\mu}$ are independent and make the natural exponential family assumption standard in BAI literature (Kaufmann et al., 2016a):
@@
For two different expectations $\mu$ and $\mu^{\prime}$ with the same exponential family, we use $d\left(\mu,\mu^{\prime}\right)$ to denote the KL-divergence from $\nu_{\mu}$ to $\nu_{\mu^{\prime}}$. Note that Assumption 1 is very general and includes a large class of distributions such as Gaussian (with known variance), Bernoulli, and Poisson distributions, where the exact parameterization is in sec. B of the appendix. Unique to this work, we assume that each arm has a cost with distribution $\boldsymbol{\nu}_{c}\coloneqq\{\nu_{c_{1}},\ldots,\nu_{c_{K}}\}$ and expectations $\boldsymbol{c}\coloneqq\{c_{1},\ldots,c_{K}\}$. We assume they satisfy the positivity assumption, which is natural in our motivating examples in real world such as ad placement or clinical trials, where the cost of each action is always bounded and all actions are not free.
@@
For any policy $\pi$, it consists of three components: (1) a sampling rule $(A_{t})_{t\geq 1}$ to select arms to interact at each round; (2) a stopping time $\tau_{\delta}$ which terminates the interaction; and (3) an arm decision rule $\widehat{a}$ to identify the best arm. As a convention of BAI with fixed confidence, we require our policy $\pi$ to be $\delta$-PAC (Probably Approximately Correct) (Kaufmann et al., 2016a), which means the algorithm should terminate in finite time and the probability of choosing the wrong best arm should be lower than the confidence level $\delta$. The definition of $\delta$-PAC is as follows:
@@ 3 Lower Bound
The proof of Theorem 1 is deferred to the appendix but primarily relies on the “transportation” lemma proposed in (Kaufmann et al., 2016b), which characterizes the theoretical hardness to distinguish the bandit model $\boldsymbol{\mu}$ from any other models $\boldsymbol{\lambda}$ where $a^{*}(\boldsymbol{\lambda})\neq a^{*}(\boldsymbol{\mu})$. Theorem 1 suggests that $\mathcal{O}(\log(1/\delta))$ cumulative cost is inevitable to identify the optimal arm, and it also characterizes the asymptotic lower bound constant $T^{*}(\boldsymbol{\mu})$.
Instance Dependent Constant $T^{*}(\boldsymbol{\mu})$: The instance dependent constant $T^{*}(\boldsymbol{\mu})$ obtained in our Theorem. 1 is different from classic best arm identification lower bounds, e.g., $T^{*}(\boldsymbol{\mu})$ in Theorem 1 of (Garivier & Kaufmann, 2016). Even though it captures the hardness of this instance in terms of the cumulative cost, $T^{*}(\boldsymbol{\mu})$ is still a vague notion in the sense that the relationship between the theoretical cumulative cost $J(\tau_{\delta})$ and model parameters, $\boldsymbol{\mu}$, $\boldsymbol{c}$, is still unclear. To better understand this mysterious constant $T^{*}(\boldsymbol{\mu})$, we present Theorem. 1 in the simple case of $2$ armed Gaussian bandits with unit variance, where $T^{*}(\boldsymbol{\mu})$ has a closed-form expression.
@@
This differs from (Garivier & Kaufmann, 2016), where $w_{a}$ is the proportion of rounds that arm $a$ is pulled. Like $T^{*}(\boldsymbol{\mu})$, there is no closed-form expression for $\boldsymbol{w}^{*}$ in general bandit models with $K\geq 3$. In the appendix, we show that one can compute the desired quantities such as $T^{*}(\boldsymbol{\mu})$ and $\boldsymbol{w}^{*}$ by similarly solving $K$ continuous equations to (Garivier & Kaufmann, 2016). Therefore, we can readily apply iterative methods such as bisection to compute these values. We summarize this procedure in the $\mathsf{ComputeProportions}$ Algorithm (Algorithm. 3 in the Appendix), which will be called regularly as a sub-routine in our proposed algorithms (Algorithm. 1).
@@ 4 Asymptotically Cost Optimal Algorithm
The cost bandit setting also encourages the algorithm to stop as early as possible, so the same stopping rules from traditional BAI (Garivier & Kaufmann, 2016) can be used. A more refined threshold can be found in (Kaufmann & Koolen, 2021). However, we will use the threshold in Algorithm 1 for the rest of the paper for simplicity. Our Proposition. 1 combines Theorem 10 and Proposition 11 from (Garivier & Kaufmann, 2016), and the proof will be provided in the appendix.
@@ 5 Low Complexity Algorithm
[…]
Proposition 2 ($\delta$-PAC).
[…]
Theorem 3.
The proof of Theorem. 3 will be delayed to the appendix. The key to the proof is to show that under our sampling rule balanced by $\sqrt{\widehat{c_{a}}(t)}$ for each arm, the empirical cost proportion $\widehat{\boldsymbol{w}}(t)$ converges to the optimal proportion $\widehat{\boldsymbol{w}}^{*}$. Then, we can apply a similar argument as the weaker version of Theorem 2 to prove the upper bound. Comparing it to Corollary. 1, we show our low-complexity algorithm is optimal in this setting. It is an important observation that in the homogeneous cost case, this algorithm reduces to a racing algorithm. It is well known that racing algorithms cannot be optimal on a general MAB model. Therefore, we cannot obtain any general optimality results. However, we will show that it enjoys surprisingly good empirical performance over a wide range of bandit models with multiple arms in the next section.
6 Numerical Experiments
Refer to caption Refer to caption (a) Arm pull trajectories
Refer to caption Refer to caption (b) Total costs.
[…]
$\mathsf{CO}$ $\mathsf{CTAS}$ $\mathsf{TAS}$ $\mathsf{d}\text{-}\mathsf{LUCB}$
Gaussian 85 1712 2410 82
Bernoulli 58 1995 2780 60
Poisson 96 3260 4633 101
@@ 7 Conclusion
In this work, we introduced a new MAB problem: Cost-Aware Best Arm Identification. We provided a new lower bound and an asymptotically optimal cost-adapted BAI algorithm. Finally, we introduced a low-complexity algorithm with promising empirical results. As a future direction, it may be interesting to explore how this algorithm can be adapted to the regret setting in either the cost adapted setting (Sinha et al., 2020), or as an ETC algorithm for carefully chosen costs.
@@ A.1 Best Arm Identification
Traditional best arm identification (BAI) has been studied for decades Bechhofer (1958). In this setting, there are two possible objectives. First, we have the fixed confidence setting, where the agent aims to minimize the sample time (or the number of samples) while ensuring the best arm is identified with probability greater than $1-\delta$. Second, there is the fixed budget setting, where the agent has a fixed number of samples and aims to minimize the probability of error.
@@ A.1.1 Fixed Confidence
In Kaufmann et al. (2016a), they introduce a non-asymptotic lower bound for BAI. Subsequently, they propose the Track and Stop algorithm ($\mathsf{TAS}$) that achieves this lower bound asymptotically. However, this algorithm requires the computation of the optimal “proportion” of arm pulls, $w^{*}$, at each iteration and requires forced exploration at the order of $\sqrt{\tau}$. The $\mathsf{TAS}$ algorithm has since been extended to a variety of other settings Jourdan et al. (2023); Garivier & Kaufmann (2021); Kato & Ariu (2021). In an attempt to ease computation time, Ménard (2019) implements a gradient ascent algorithm to get around having to compute the optimal proportions at each iteration.
@@
Prior to these “model-based” methods, there were many “confidence-based” algorithms Kalyanakrishnan et al. (2012); Kaufmann & Kalyanakrishnan (2013); Gabillon et al. (2012) focused on constructing high-probability confidence intervals. The stopping time is then when one confidence interval was disjoint from and greater than all the rest. While these algorithms are more computationally efficient than the aforementioned “model-based” algorithms, they do not perform nearly as well. These algorithms ideas were heavily inspired by the regret minimization setting and the heralded UCB algorithm Lai & Robbins (1985); Cappé et al. (2013); Lai (1987).

### v2

@@ Abstract
In this paper, we study a best arm identification problem with dual objects. In addition to the classic reward, each arm is associated with a cost distribution and the goal is to identify the largest reward arm using the minimum expected cost. We call it Cost Aware Best Arm Identification (CABAI), which captures the separation of testing and implementation phases in product development pipelines and models the objective shift between phases, i.e., cost for testing and reward for implementation. We first derive a theoretical lower bound for CABAI and propose an algorithm called $\mathsf{CTAS}$ to match it asymptotically. To reduce the computation of $\mathsf{CTAS}$, we further propose a simple algorithm called Chernoff Overlap (CO), based on a square-root rule, which we prove is optimal in simplified two-armed models and generalizes well in numerical experiments. Our results show that (i) ignoring the heterogeneous action cost results in sub-optimality in practice, and (ii) simple algorithms can deliver near-optimal performance over a wide range of problems.
@@ 1 Introduction
The stochastic multi-armed bandit (MAB) (Thompson, 1933; Robbins, 1952) is a classic model which has widespread applications, from content recommendation (Kohli et al., 2013), resource allocation (Liu et al., 2020), clinical trials (Villar et al., 2015), to efficient ad placement. A multi-armed bandit problem involves an agent and an environment, which is represented by a set of actions (arms) with distinct underlying reward distributions. At each round, the agent will choose one of the arms and then obtain a random reward generated from the associated distribution. Most existing studies formulate MAB either as the best arm identification (BAI) problem (Kaufmann et al., 2016; Garivier & Kaufmann, 2016) where the agent intends to identify the highest reward arm as quickly as possible or as a regret minimization problem (Auer, 2002; Garivier & Cappé, 2011) where the goal is to maximize the cumulative reward over a time-horizon. Both formulations have been well-developed and successful in balancing the trade-off between exploration and exploitation.
@@
Algorithm Optimal? $w_{1}(t)$ $(1.5,1)$ $w_{2}(t)$ $(1,0.1)$ $w_{3}(t)$ $(0.5,0.01)$
$\mathsf{TAS}$$\times$ $0.46$ $0.46$ $0.08$
$\mathsf{CTAS}$ $0.23$ $0.72$ $0.05$
@@ 1.1 Related Work
BAI with Fixed Confidence: BAI has been studied for many years and was originally proposed in Bechhofer (1958). In this paper, we consider a subset known as the fixed confidence setting, where the agent aims to minimize the sample complexity while ensuring the best arm is identified with probability at least $1-\delta$. Here, $\delta$ is a pre-specified confidence level, and such algorithms are called $\delta$-PAC. In Kaufmann et al. (2016), the authors introduce a non-asymptotic lower bound for this setting. Subsequently, they propose the Track and Stop algorithm ($\mathsf{TAS}$) that matches this lower bound asymptotically. The $\mathsf{TAS}$ algorithm has since been extended to various other settings (Jourdan et al., 2023; Garivier & Kaufmann, 2021; Kato & Ariu, 2021). Before it, researchers proposed “confidence-based” algorithms, e.g., $\mathsf{KL}$-$\mathsf{LUCB}$ (Kaufmann & Kalyanakrishnan, 2013), $\mathsf{UGapE}$ (Gabillon et al., 2012), which rely on constructing high-probability confidence intervals. They are more computationally feasible than $\mathsf{TAS}$ inspired algorithms, but with few theoretical guarantees.
BAI with Safety Constraints: A formulation similar to our paper is BAI with safety constraints (Wang et al., 2022). As a motivating example, they consider the clinical trial setting, where each drug is associated with a dosage and the dosage has an associated safety level. They attempt to identify the best drug and dosage for fixed confidence without violating the safety level. Similarly, Hou et al. (2023) attempts to identify the best arm subject to a constraint on the variance. Our formulation is distinct from them because the agent is free to perform any action. In Chen et al. (2022b; a), they formulate safety constraints as a constrained optimization problem. They explore and show that allowing minimal constraint violations can provide significant improvement in the regret setting. This is distinct from the BAI setting explored in this paper.
Multi-fidelity BAI: An alternative formulation that considers cost is the multi-fidelity formulation introduced in Kandasamy et al. (2016) and recently considered in the best arm identification regime (Poiani et al., 2022; Wang et al., 2023). In this setting, along with choosing an arm, the agent chooses the desired fidelity or “level of accuracy” of the mean estimate. Each fidelity incurs a cost, where higher fidelity incurs a larger cost but provides more accurate estimate. This setting clearly differs from ours because the cost of each fidelity is known a priori and is controllable through choice of fidelity.
@@ 2 Preliminaries
We study a model similar to the fixed-confidence BAI in stochastic $K$-armed bandits. We denote the set of arms as $\mathcal{A}\coloneqq\{1,2,\cdots,K\}$. Each arm $a$ is associated with a reward distribution $\boldsymbol{\nu}_{\mu}=\{\nu_{\mu_{1}},\ldots,\nu_{\mu_{K}}\}$ with expectations $\boldsymbol{\mu}\coloneqq\{\mu_{1},\mu_{2},\cdots,\mu_{K}\}$. We assume $\boldsymbol{\nu}_{\mu}$ are independent and make the natural exponential family assumption standard in BAI literature (Kaufmann et al., 2016):
@@
For two different expectations $\mu$ and $\mu^{\prime}$ with the same exponential family, we use $d\left(\mu,\mu^{\prime}\right)$ to denote the KL-divergence from $\nu_{\mu}$ to $\nu_{\mu^{\prime}}$. Note that Assumption 1 is very general and includes a large class of distributions such as Gaussian (with known variance), Bernoulli, and Poisson distributions by considering the following choice of parameters:
Unique to this work, we assume that each arm has a cost with distribution $\boldsymbol{\nu}_{c}\coloneqq\{\nu_{c_{1}},\ldots,\nu_{c_{K}}\}$ and expectations $\boldsymbol{c}\coloneqq\{c_{1},\ldots,c_{K}\}$. We assume they satisfy the positivity assumption, which is natural in our motivating examples in real world such as ad placement or clinical trials, where the cost of each action is always bounded and all actions are not free.
@@
For any policy $\pi$, it consists of three components: (1) a sampling rule $(A_{t})_{t\geq 1}$ to select arms to interact at each round; (2) a stopping time $\tau_{\delta}$ which terminates the interaction; and (3) an arm decision rule $\widehat{a}$ to identify the best arm. As a convention of BAI with fixed confidence, we require our policy $\pi$ to be $\delta$-PAC (Probably Approximately Correct) (Kaufmann et al., 2016), which means the algorithm should terminate in finite time and the probability of choosing the wrong best arm should be lower than the confidence level $\delta$. The definition of $\delta$-PAC is as follows:
@@ 3 Lower Bound
The proof of Theorem 1 is deferred to the appendix but primarily relies on the “transportation” lemma proposed in Kaufmann et al. (2016), which characterizes the theoretical hardness to distinguish the bandit model $\boldsymbol{\mu}$ from any other models $\boldsymbol{\lambda}$ where $a^{*}(\boldsymbol{\lambda})\neq a^{*}(\boldsymbol{\mu})$. Theorem 1 suggests that $\mathcal{O}(\log(1/\delta))$ cumulative cost is inevitable to identify the optimal arm, and it also characterizes the asymptotic lower bound constant $T^{*}(\boldsymbol{\mu})$.
Instance Dependent Constant $T^{*}(\boldsymbol{\mu})$: The instance dependent constant $T^{*}(\boldsymbol{\mu})$ obtained in our Theorem. 1 is different from classic best arm identification lower bounds, e.g., $T^{*}(\boldsymbol{\mu})$ in Theorem 1 of Garivier & Kaufmann (2016). Even though it captures the hardness of this instance in terms of the cumulative cost, $T^{*}(\boldsymbol{\mu})$ is still a vague notion in the sense that the relationship between the theoretical cumulative cost $J(\tau_{\delta})$ and model parameters, $\boldsymbol{\mu}$, $\boldsymbol{c}$, is still unclear. To better understand this mysterious constant $T^{*}(\boldsymbol{\mu})$, we present Theorem. 1 in the simple case of $2$ armed Gaussian bandits with unit variance, where $T^{*}(\boldsymbol{\mu})$ has a closed-form expression.
@@
This differs from Garivier & Kaufmann (2016), where $w_{a}$ is the proportion of rounds that arm $a$ is pulled. Like $T^{*}(\boldsymbol{\mu})$, there is no closed-form expression for $\boldsymbol{w}^{*}$ in general bandit models with $K\geq 3$. In the appendix, we show that one can compute the desired quantities such as $T^{*}(\boldsymbol{\mu})$ and $\boldsymbol{w}^{*}$ by similarly solving $K$ continuous equations to Garivier & Kaufmann (2016). Therefore, we can readily apply iterative methods such as bisection to compute these values. We summarize this procedure in the $\mathsf{ComputeProportions}$ Algorithm (Algorithm. 3 in the Appendix), which will be called regularly as a sub-routine in our proposed algorithms (Algorithm. 1).
@@ 4 Asymptotically Cost Optimal Algorithm
Remark 1.
For Theorem 5 and the almost sure optimality result, we do not need such a fast rate of convergence and only require that for every $a$, $N_{a}(t)\to\infty$.
@@
The cost bandit setting also encourages the algorithm to stop as early as possible, so the same stopping rules from traditional BAI (Garivier & Kaufmann, 2016) can be used. A more refined threshold can be found in Kaufmann & Koolen (2021). However, we will use the threshold in Algorithm 1 for the rest of the paper for simplicity. Our Proposition. 1 combines Theorem 10 and Proposition 11 from Garivier & Kaufmann (2016), and the proof will be provided in the appendix.
@@ 5 Low Complexity Algorithm
[…]
Proposition 2 ($\delta$-PAC).
[…]
Theorem 3.
The proof of Theorem. 3 will be delayed to the appendix. The key to the proof is to show that under our sampling rule balanced by $\sqrt{\widehat{c_{a}}(t)}$ for each arm, the empirical cost proportion $\widehat{\boldsymbol{w}}(t)$ converges to the optimal proportion $\widehat{\boldsymbol{w}}^{*}$. Then, we can apply a similar argument as the weaker version of Theorem 2 to prove the upper bound. Comparing it to Corollary. 1, we show our low-complexity algorithm is optimal in this setting. It is an important observation that in the homogeneous cost case, this algorithm reduces to a racing algorithm. It is well known that racing algorithms cannot be optimal on a general MAB model. However, we will show that it enjoys surprisingly good empirical performance over a wide range of bandit models with multiple arms in the next section. Establishing a provable suboptimality gap is an interesting future research problem.
6 Numerical Experiments
$\mathsf{CO}$ $\mathsf{CTAS}$ $\mathsf{TAS}$ $\mathsf{d}\text{-}\mathsf{LUCB}$
Gaussian 85 1712 2410 82
Bernoulli 58 1995 2780 60
Poisson 96 3260 4633 101
@@ 7 Conclusion
In this work, we introduced a new MAB problem: Cost-Aware Best Arm Identification. We provided a new lower bound and an asymptotically optimal cost-adapted BAI algorithm. Finally, we introduced a low-complexity algorithm with promising empirical results. As a future direction, it may be interesting to explore how this algorithm can be adapted to the regret setting in either the cost adapted setting (Sinha et al., 2021), or as an ETC algorithm for carefully chosen costs.
Acknowledgments
This work is supported in part by NSF under grants 2112471, 2134081, 2207548, 2240981, and 2331780.
(a) Arm pull trajectories
(b) Total costs.
[…]
@@ A.1 Best Arm Identification
Traditional best arm identification (BAI) has been studied for decades (Bechhofer, 1958). In this setting, there are two possible objectives. First, we have the fixed confidence setting, where the agent aims to minimize the sample time (or the number of samples) while ensuring the best arm is identified with probability greater than $1-\delta$. Second, there is the fixed budget setting, where the agent has a fixed number of samples and aims to minimize the probability of error.
@@ A.1.1 Fixed Confidence
In Kaufmann et al. (2016), they introduce a non-asymptotic lower bound for BAI. Subsequently, they propose the Track and Stop algorithm ($\mathsf{TAS}$) that achieves this lower bound asymptotically. However, this algorithm requires the computation of the optimal “proportion” of arm pulls, $w^{*}$, at each iteration and requires forced exploration at the order of $\sqrt{\tau}$. The $\mathsf{TAS}$ algorithm has since been extended to a variety of other settings (Jourdan et al., 2023; Garivier & Kaufmann, 2021; Kato & Ariu, 2021). In an attempt to ease computation time, Ménard (2019) implements a gradient ascent algorithm to get around having to compute the optimal proportions at each iteration.
@@
Prior to these “model-based” methods, there were many “confidence-based” algorithms (Kalyanakrishnan et al., 2012; Kaufmann & Kalyanakrishnan, 2013; Gabillon et al., 2012) focused on constructing high-probability confidence intervals. The stopping time is then when one confidence interval was disjoint from and greater than all the rest. While these algorithms are more computationally efficient than the aforementioned “model-based” algorithms, they do not perform nearly as well. These algorithms ideas were heavily inspired by the regret minimization setting and the heralded UCB algorithm (Lai & Robbins, 1985; Cappé et al., 2013; Lai, 1987).